OpenStudy (anonymous):

convert to rectangular form r=6csc theta

OpenStudy (sidsiddhartha):

use y=rsin(theta) and x=rcos(theta) now tan(theta)=y/x so calculate cosec(theta) its=csc(theta) = sqrt(x^2+y^2)/y and r=sqrt(x^2+y^2) now just substitute

OpenStudy (asnaseer):

@ddd231 - do you know what csc(theta) is in terms of the standard sin/cos/tan functions?

OpenStudy (anonymous):

yes, and is the answer y=6?

OpenStudy (asnaseer):

correct

OpenStudy (sidsiddhartha):

yup very good

OpenStudy (anonymous):

is there no need for x to be shown in the answer?

OpenStudy (asnaseer):

so your equation can be written as:\[r=6\csc(\theta)=6/\sin(\theta)\]if you multiply both sides by \(\sin(\theta)\) you should get to the desired answer very quickly

OpenStudy (asnaseer):

there is no x in the answer as this equation represents the line y=6

OpenStudy (anonymous):

oh ok thank you both

OpenStudy (asnaseer):

np :)

OpenStudy (anonymous):

If both of you all still have time I want to check my work with another problem. @asnaseer @sidsiddhartha

OpenStudy (anonymous):

This is another conversion from polar to rectangular form. r=4/(1+2sin(theta)) would that become 3y^2=x^2?

OpenStudy (asnaseer):

@ddd231 - please post each question separately. That way each question is self contained and more focused towards a particular question. So can you please close this question and post the new one in the list to the left. Thanks.

OpenStudy (anonymous):

ah, alright sorry